OpenStudy (flvskidd):

You are having a conference call with the CEO of a paper company. You have interpreted the number of trees cut down versus profit as the function P(x) = –x4 + x3 + 7x2 − x − 6. Describe to the CEO what the graph looks like and, in general, how to sketch the graph without using technology. Use complete sentences, and focus on the end behaviors of the graph and where the company will break even (where P(x) = 0).

OpenStudy (anonymous):

Okay. We can factor P(x) using division. \[P(x) = (x - 3)(-x^3 - 2x^2 + x + 2)\] Using division once more, \[P(x) = (x - 3)(x + 2)(-x^2 + 1)\] Set f(x) equal to zero. \[0 = (x - 3)(x + 2)(-x^2 + 1)\] Set the factors equal to zero. \[x - 3 = 0 \] \[ x + 2 = 0\] \[-x^2 + 1 = 0\] \[x = 3 \] \[ x = -2\] \[-x^2 = -1\] \[x^2 = 1\] \[ x = -1\] \[and\] \[x = 1\] Now that we have all the info we need, lets describe the graph. This is polynomial with degree 4. Since the coefficient of the leading term is negative and the degree is positive, P(x) starts off to be increasing, and will end decreasing. In addition, the multiplicity of each factor is odd, which means the roots do not repeat, so P(x) intercepts the x-axis four times. The first two factors act like a linear function with positive slopes. The other factor acts like a negative parabola. The y-intercept is (0, -6). Using this information as best I have described, can the CEO graph P(x) by hand?

OpenStudy (anonymous):

Now that is your part to figure out :)

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @FLVSKidd